OpenStudy (anonymous):

Find the length of the hypotenuse. Show all work. (3pts)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[hyp= \sqrt{x^2+y^2}\] \[hyp = \sqrt{(\sqrt3)^2+(2\sqrt3)^2}\] \[hyp = \sqrt{3+4(3)}\] \[hyp =\sqrt{15}\]

OpenStudy (anonymous):

thank you ! so the sqrt of 15 squared is just 15 ?

OpenStudy (anonymous):